# Disentangled and Self-Explainable Node Representation Learning

### v1

@@ A.4 Results for synthetic datasets
In Figure A4 we plot results for the plausibility metric on link prediction and node classification, on the top and the bottom respectively. Plausibility seems to benefit larger dimension values for DiSe methods and DW+Dine for link prediction. Figure A5 shows the corresponding downstream task accuracy results.
@@
Figure A4: Plausibility results on synthetic datasets with varying feature dimensions size.
@@
Figure A5: Downstream tasks results on synthetic datasets with varying feature dimensions size.
@@
Given an instance (can be a node embedding $\mathbf{h}(v)$ or a edge embedding $\mathbf{h}(u,v)$) we compute feature importance values with SHAP $\{\Psi^{(\mathcal{V})}_{j}\}_{j=1...K}$ or $\{\Psi^{(\mathcal{E})}_{j}\}_{j=1...K}$ and the corresponding task-based graph masks:
$\mathbf{B}^{(j)}(\mathcal{V})\in\mathbb{R}_{\geq 0}^{V};~{}~{}~{}B^{(j)}(v;b)=\mathrm{max}\{0,\Psi^{(\mathcal{V})}_{j}(v;b)\}$ (10)
$\mathbf{B}^{(j)}(\mathcal{E})\in\mathbb{R}_{\geq 0}^{E};~{}~{}~{}B^{(j)}(u,v;b)=\mathrm{max}\{0,\Psi^{(\mathcal{E})}_{j}(u,v;b)\}$ (11)

### v2

@@ Appendix F Explanations for Synthetic Datasets
Figure A11: Plausibility results on synthetic datasets (link prediction on the top panel, binary node classification on the bottom panel) with varying feature dimensions size.
@@
Figure A12: Downstream tasks results on synthetic datasets (link prediction on the top panel, binary node classification on the bottom panel) with varying feature dimensions size.
Appendix G Influence of Entropy Regularizer and Model Depth for Wiki Dataset
Figure A8 explores how both the entropy-regularization coefficient, $\lambda_{ent}$, and the GCN depth in DiSe-GAE influence the interpretability metrics on the Wiki dataset. Increasing $\lambda_{ent}$ and using deeper networks boost Topological Alignment and Overlap Consistency, but it also reduce Sparsity, revealing a trade-off between compactness and explanatory power. Positional Coherence is weakly sensitive to $\lambda_{ent}$, and tends to improve with shallower architectures.
Appendix H Qualitative Examples for Cora Dataset
Tables A7 and A8 presents qualitative examples regarding the embedding interpretability on the Cora citation graph. The dataset comprises seven classes, each corresponding to a distinct computer science topic. We employ different embeddings as predictor variables for the node classification task with a one-vs-rest multiclass strategy using Logistic Regression. Comparing DiSe-GAE with the vanilla GAE and GAE+DINE, we selected for each class the most task-relevant embedding dimension corresponding to the highest coefficient of the linear classifier. For every selected dimension, we analyze different properties of the corresponding anchor subgraph, computed with Eq. (2). First, without inspecting node class labels, the Topological Alignment scores and the Density scores (i.e., the ratio of the number of edges with respect to the maximum possible edges in the subgraph) tell us that DiSe-GAE produces denser subgraphs compared to the overall graph, indicating stronger structural cohesion in the most task-relevant subgraph. Moreover, considering the bar plots that display distributions of class labels inside the subgraphs, nodes grouped within a single dimension of DiSe-GAE share research themes more consistently. This is quantified in the lower entropy of labels’ distributions within each subgraph. Furthermore, by looking at the top-20 most frequent keywords extracted from each cluster (after lemmatization and stop-words removal), we represent the word-cloud for each dimension, indicating clusters that are more thematically coherent and easier to interpret. For example, looking at the most predictive dimensions for the label Probabilistic methods, we observe with DiSe-GAE we identify papers mentioning topic-relevant keywords such as "bayesian netwoks", "inference" and "probabilistic"; while papers identified with GAE+DINE frequently refer to "reinforcement learning" and "control". These qualitative examples therefore show that our model learns latent dimensions whose structural and semantic meanings are markedly more human-intelligible than those of the competing methods.
Table A7: Most task-relevant dimensions in different methods for node labels Neural Networks, Reinforcement Learning, and Probabilistic Methods and their interpretations based on topic and text information in Cora.
Node Label DiSE-GAE (Macro-F1=.864) GraphAE (Macro-F1=.865) GAE+DINE (Macro-F1=.838)
Topolog. Align.=.874 Topolog. Align.=.272 Topolog. Align.=.469
@@ Appendix H Qualitative Examples for Cora Dataset
Labels Entropy=.216 Labels Entropy=.876 Labels Entropy=.355
Neural Networks [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Topolog. Align.=.987 Topolog. Align.=.579 Topolog. Align.=.871
@@
Labels Entropy=.151 Labels Entropy=.832 Labels Entropy=.494
Reinforcement Learning [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Topolog. Align.=.897 Topolog. Align.=.232 Topolog. Align.=.720
[…]
Labels Entropy=.242 Labels Entropy=.699 Labels Entropy=.619
Probabilistic Methods [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table A8: Most task-relevant dimensions in different methods for node labels Theory, Genetic Algorithms, and Rule Learning and their interpretations based on topic and text information in Cora.
Node Label DiSE-GAE (Macro-F1=.864) GraphAE (Macro-F1=.865) GAE+DINE (Macro-F1=.838)
Topolog. Align.=.937 Topolog. Align.=.402 Topolog. Align.=.676
[…]
Labels Entropy=.417 Labels Entropy=.878 Labels Entropy=.349
Theory [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Topolog. Align.=.992 Topolog. Align.=.309 Topolog. Align.=.608
[…]
Labels Entropy=.050 Labels Entropy=.755 Labels Entropy=.434
Genetic Algorithms [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Topolog. Align.=.693 Topolog. Align.=.594 Topolog. Align.=.362
[…]
Labels Entropy=.516 Labels Entropy=.805 Labels Entropy=.647
Rule Learning [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]